OpenStudy (anonymous):

find dy/dx if xcos(y) - y^2 = 4 - x^2

OpenStudy (anonymous):

i got - [2x] / [ xsin(y) - cos(y) ]

OpenStudy (cwrw238):

implicit differentiation: cos(y) + x* -sin(y) * dy/dx - 2y* dy/dx = -2x 2y* dy/dx - sin(y) x dy/dx = -2x - cos(y) dy/dx = (-2x - cos(y) ) / (2y - x sin(y) )

OpenStudy (cwrw238):

hmm - i'm just checking my work....

OpenStudy (cwrw238):

i think i'm right - i'll check it out on wolfram

OpenStudy (cwrw238):

its 2x + cos (y) ---------- x sin(y) - 2y which is same as my answer except they have multiplied top and bottom by -1 to make the whole thing positive

OpenStudy (anonymous):

what's the first step.. i don't think i get it.

myininaya (myininaya):

\[(xcos(y))'=(x)'\cos(y)+x(\cos(y))'=1 \cos(y)+x y' (-\sin(y))=\cos(y)-xy'\sin(y)\] \[(y^2)'=2yy'\] \[(4-x^2)=(4)'-(x^2)'=0-2x=-2x\] So we have \[\cos(y)-xy'\sin(y)-2yy'=-2x\]

OpenStudy (anonymous):

what's deriv of cos(y)

OpenStudy (anonymous):

-sin(y) * y' right?

OpenStudy (cwrw238):

differentiating x cos(y) is done using the product rule as myin has done above using the y' notation cos)y) when differentiating w r t to x you regard y as a function of f x an use chain rule

OpenStudy (anonymous):

yes i know to use chain

OpenStudy (cwrw238):

yes -siny * y' is correct

OpenStudy (cwrw238):

ok- sorry

OpenStudy (anonymous):

so i get -xy'sin(y) + cos(y) = -2x

OpenStudy (anonymous):

is that correct so far?

OpenStudy (cwrw238):

no you forgot derivative of y^2 which is 2y* y'

OpenStudy (anonymous):

ahhhhhhhhhhhhhh

OpenStudy (cwrw238):

-xy'sin(y) + cos(y) - 2y y' = -2x

OpenStudy (cwrw238):

i just saw an error - the second line should start with -2

OpenStudy (cwrw238):

so the denominator of the answer should read x sin(y) + 2y

OpenStudy (cwrw238):

phew!!!

OpenStudy (anonymous):

-2y'xsin(y)-2y = -2x

OpenStudy (anonymous):

good till there?

OpenStudy (cwrw238):

yea

OpenStudy (cwrw238):

-2y'xsin(y)-2y y' = -2x

OpenStudy (anonymous):

so i get: - [2x] / [2xsin(y)+2y]

OpenStudy (anonymous):

err

OpenStudy (cwrw238):

hmm - i'm going crazy here ( not your fault) - lets have another check

OpenStudy (cwrw238):

you've missed out the cos(y) alien

OpenStudy (anonymous):

nah i forgot cos... lemme do it again

OpenStudy (anonymous):

(-xy'sin(y))+(cos(y))-[2y*y'] = -2x

OpenStudy (anonymous):

so far so good?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

ok so how do i combine the y'

OpenStudy (cwrw238):

take out the y' form 2 terms containing it and subtract cos(y) from both sides

OpenStudy (cwrw238):

y'( -x sin(y) - 2y) = -2x - cos(y)

OpenStudy (anonymous):

so finally i get: - [2x + cos(y)] / [xsin(y) + 2y]

OpenStudy (cwrw238):

no minus in the numerator

OpenStudy (anonymous):

i factored out the minus from the whole thing

OpenStudy (anonymous):

so before i did that it read: [-2x-cos(y)] / [xsin(y)+2y]

OpenStudy (anonymous):

err

OpenStudy (anonymous):

i mean that [-2x-cos(y)] / [-xsin(y)+2y]

OpenStudy (anonymous):

ack

OpenStudy (anonymous):

i mean this for real this time: [-2x-cos(y)] / [-xsin(y)-2y]

OpenStudy (cwrw238):

itd -2x - cos(y) / -xsin(y) - 2y - yes so when you multiply top and bottom by -1 all signs become positive

OpenStudy (anonymous):

hurray

OpenStudy (cwrw238):

yayyy!!

OpenStudy (anonymous):

i quit math now.

OpenStudy (cwrw238):

lol